# SparseVLM: Visual Token Sparsification for Efficient Vision-Language Model Inference

### v2

@@ Abstract
In vision-language models (VLMs), visual tokens usually consume a significant amount of computational overhead, despite their sparser information density compared to text tokens. To address this, most existing methods learn a network to prune redundant visual tokens and require additional training data. Differently, we propose an efficient training-free token optimization mechanism dubbed SparseVLM without extra parameters or fine-tuning costs. Concretely, given that visual tokens complement text tokens in VLMs for linguistic reasoning, we select visual-relevant text tokens to rate the significance of vision tokens within the self-attention matrix extracted from the VLMs. Then we progressively prune irrelevant tokens. To maximize sparsity while retaining essential information, we introduce a rank-based strategy to adaptively determine the sparsification ratio for each layer, alongside a token recycling method that compresses pruned tokens into more compact representations. Experimental results show that our SparseVLM improves the efficiency of various VLMs across a range of image and video understanding tasks. In particular, LLaVA equipped with SparseVLM reduces 61% $\sim$ 67% FLOPs with a compression ratio of 78% while maintaining 93% of the accuracy. Our code is available at https://github.com/Gumpest/SparseVLMs.
@@ 1 Introduction
Benefiting from tremendous advancements in large language models (LLMs) (Radford et al., 2019; Brown et al., 2020; Achiam et al., 2023; Touvron et al., 2023; Peng et al., 2023; Bi et al., 2024), the realm of vision-language models (VLMs) has undergone a revolutionary progress. To combine visual signals with textual semantics, the mainstream practice in VLMs (Team et al., 2023; Bai et al., 2023b; Chen et al., 2023; Li et al., 2024b; 2023a) employs sequential visual representation, where images are extracted into vision tokens and sent into an LLM decoder. With modal alignment and instruction fine-tuning (Du et al., 2021; Liu et al., 2023a; Zhu et al., 2023b), recent VLMs successfully adapt LLMs to the vision domain and inherit their perception and reasoning abilities.
Despite the promising performance, further incorporation of visual tokens inevitably introduces a huge memory and computation overhead when compared to LLMs, particularly for high-resolution images (Li et al., 2024b) and multi-frame videos (Lin et al., 2023). For instance, a $672\times 672$ image in LLaVA (Liu et al., 2024) yields $2304$ vision tokens that span over half of the context length. However, the information in images is typically more sparse than in natural languages (Marr, 2010), resulting in inefficiency in directly processing both modalities. To address this, existing methods extract more compact image representations by modifying the image encoder or projector (Alayrac et al., 2022; Li et al., 2024a; Dai et al., 2023; Cha et al., 2024). While some recent works further sparsify vision tokens during the decoding (Ye et al., 2024; Chen et al., 2024b; Cai et al., 2024), they still ignore the guidance from the language tokens, which contradicts the multimodality paradigm. We argue that visual tokens should be sparsified adaptively based on the question prompt, as the model might focus on different parts (e.g., foreground or background) when dealing with various questions, as shown in Figure 1. Furthermore, current approaches generally train a network to prune redundant visual tokens and require additional training data (Li et al., 2024a; Ye et al., 2024).
In this paper, we introduce a text-guided training-free framework dubbed SparseVLM for efficient vision language model inference. We reuse the self-attention matrix of visual-text tokens directly from the decoder layers without extra training parameters for sparsification. We ascertain that not all prompt tokens should be considered as some could be less relevant, which leads to inaccurate correlation results and downgrades the performance of sparse inference. Specifically, our SparseVLM first identifies text tokens strongly correlated with visual signals via cross-attention. Then, we measure the contribution of visual tokens to the selected visual-relevant text tokens (raters) and adaptively prune the insignificant vision tokens. Instead of directly discarding the pruned tokens, we further recycle and cluster them to reconstruct more compact tokens to minimize information loss. Due to the information density varying for different image inputs, we employ the rank of the attention matrix to indicate the redundancy level and set an adaptive sparsification ratio accordingly.
The proposed method is simple yet practical. It can act as a plug-and-play module to improve the efficiency of VLMs without additional fine-tuning. Extensive experiments demonstrate that our SparseVLM effectively reduces the computation of various VLMs without sacrificing their performance in a wide range of image and video understanding tasks. For example, LLaVA (Liu et al., 2024) equipped with SparseVLM achieves a 4.5×4.5$\times$4.5 × compression rate while maintaining $93\%$ of its original performance. Alternatively, the latency (CUDA time) can decrease by $53.9\%$ with only a $13\%$ drop in accuracy. To investigate the effectiveness of our method in the video tasks, we further apply SparseVLM to VideoLLaVA (Lin et al., 2023) to additionally compress temporal frames. Without complex design changes, SparseVLM can sparsify video frames into an adaptive number of vision tokens and outperforms existing vision compression methods in video question-answering benchmarks. For instance, our method average exceeds FastV (Chen et al., 2024b) by $34.4\%$.
In summary, the contributions of this paper are threefold:
  1. 1.
    We introduce a novel sparsification framework dubbed SparseVLM for vision-language models. To the best of our knowledge, it is the first attempt to explore the potential of text-aware guidance for efficient inference of VLMs, where additional training is unnecessary.
  2. 2.
    In the framework, we first assign visual-relevant text tokens as raters, to judge the importance of vision tokens. Additionally, the rank of the attention logits is employed to reflect the redundancy and adaptively prune VLMs. Finally, we recycle partial tokens from the pruned pools and reconstruct them to accommodate more information within fewer slots.
  3. 3.
    We apply our SparseVLM framework to both image and video VLMs and conduct extensive experiments across various benchmarks. Our approach consistently outperforms the existing state-of-the-art method FastV by $7.7\%\sim 14.8\%$ on LLaVA, $10.2\%\sim 21.6\%$ on MiniGemini, and $34.4\%$ on VideoLLaVA.
@@
Figure 1: Visualization of different visual token sparsification methods. Unlike previous methods with text-agnostic visual sparsification (c) e.g., recent FastV (Chen et al., 2024b), our SparseVLM (b) is guided by question prompts to select relevant visual patches. Best viewed in color.
@@ 2 Related Work
Vision-Language Models. With the impressive success of large language models (LLMs) (Achiam et al., 2023; Touvron et al., 2023; Bai et al., 2023a), recent works on generative vision-language models (VLMs) (Liu et al., 2023a; Chen et al., 2023; Li et al., 2024b) improve multimodal comprehension and generation by receiving a long visual token sequence. Moreover, processing higher-resolution images inevitably entails an exponential growth in the length of the visual sequence. For example, LLaVA encodes $336\times 336$ images into $576$ tokens (Liu et al., 2024) and processes images with a greater resolution of $672\times 672$ into $2304$ tokens (Liu et al., 2023a). Along the same lines, mini-Gemini-HD (Li et al., 2024b) converts images into $2880$ vision tokens based on the standard of high resolution $1536\times 1536$ and low resolution $672\times 672$. Comprehending videos or multiple images inherently necessitates increased token slots for visual signals. The VideoLLaVA (Lin et al., 2023) and VideoPoet (Kondratyuk et al., 2023) models allocate thousands of tokens to afford the representation of multiple frames. However, the large number of vision tokens leads to a huge bottleneck for computational infrastructure. Further research and development in sparsification technologies are urged to overcome these hurdles and fully unleash the potential of VLMs.
@@
Visual Compression for VLMs. Compression of vision tokens is necessary because, on the one hand, their quantity is usually tens to hundreds of times that of language tokens. On the other hand, visual signals are inherently redundant compared to dense human-designed texts (Marr, 2010). Past efforts to address the above problem can be categorized into two directions. The first one centers on the vision tower or projection of visual modality and cuts vision tokens with external modules. For instance, LLaMA-VID (Li et al., 2024a) exploits the use of Q-Former with context token while DeCo (Yao et al., 2024) employs an adaptive pooling to downsample the visual tokens at the patch level. The other type methods (Ye et al., 2024; Chen et al., 2024b; Cai et al., 2024) go deeper into the text modality and sparsify vision tokens during LLM decoding, but they still neglect the guidance from the text tokens. In our paper, SparseVLM takes note of this and improves performance upon it.
Token Merging for VLMs. Token merging has recently received much attention in the field of VLMs where its algorithms mainly fall into two directions. One focuses on the matching algorithm using the Bipartite Soft Matching (BSM) (Bolya et al., 2022). For example, ToMe (Bolya et al., 2022) prunes according to the relevance between visual tokens and text and merges both modalities through the BSM algorithm. Other methods rely on clustering methods for token merging. For instance, LLaVolta (Chen et al., 2024a) proposes the visual context compressor (an average pooling) to merge the output tokens from the vision tower and progressively enhance the VLMs’ efficiency by training. Inspired by TCFormer (Zeng et al., 2024), we apply $k$-nearest neighbor density peak aggregation algorithm (Rodriguez, 2014) to the field of vision-language models. Unlike TCFormer, our method focuses on merging the dropped tokens and makes full use of their detailed information.
3 Proposed Approach: SparseVLM
@@
In this section, we present our SparseVLM framework for efficient VLM inference. We first review the attention mechanism in VLMs and then introduce the detailed strategies for our visual sparsification, including visual significance estimation, relevant text token selection, and sparsification level adaptation. We further propose the token recycling method to reduce information loss and provide a theoretical analysis of computation savings. The overall architecture is shown in Figure 2.
@@
Figure 2: The architecture of SparseVLM. In stage (a), text raters are pre-selected before entering the sparsification LLM. In stage (b), adaptive sparsification is performed on LLM layers, involving computing redundancy and the recycling of reconstructed tokens. Best viewed in color.
@@ 3.1 Preliminary: Attention in VLM Decoders
The VLM decoders typically use the causal self-attention in the original transformer model (Vaswani et al., 2017) for token interactions. Without loss of generality, we discuss the condition of single-head attention. Formally, the self-attention matrix with logits ${\bm{A}}\in\mathbb{R}^{L\times L}$, where $L$ denotes the length of total tokens (e.g., system, image, and question prompt tokens), is computed by
@@
${\bm{A}}=\text{Attention}({\bm{Q}},{\bm{K}})=\text{Softmax}\left(\frac{{\bm{Q}}{\bm{K}}^{T}}{\sqrt{D}}\right),$ (1)
@@
where the scalar $D$ represents the matrix dimension, and the ${\bm{Q}}\in\mathbb{R}^{L\times D}$ and ${\bm{K}}\in\mathbb{R}^{L\times D}$ are the query and key matrices, respectively. The keys and queries in a self-attention layer are computed in parallel by using multi-layer perceptrons (MLPs) to transform the input hidden states ${\bm{H}}$ into a common space, where aligned interactions between different modalities occur.
3.2 Sparsification Guidance from Text to Visual Modality
@@
Estimation of Visual Token Significance. For a multimodal vision-language model, we should consider its impact on other modalities when deleting a single modal information. In our case, we need to understand how relevant a visual token is to the textual tokens in order to determine whether it should be removed. Therefore, we naturally come up with reusing the self-attention logits in VLMs transformer layers as a reference, since they already contain language-to-vision query results.
@@
In particular, we take the interaction between the query-dimensional part of logits of the textual modality and the key-dimensional part of the visual modality as the basis for sparsification priority matrix ${\bm{P}}\in\mathbb{R}^{L_{t}\times L_{v}}$, where $L_{t}$ and $L_{v}$ are the lengths of text and vision tokens, that is defined by
${\bm{P}}={\bm{A}}[i_{t},i_{v}],$ (2)
with
$i_{t}\in\{x|{\bm{A}}[x,:]={\mathbb{L}}\},\ \ \ i_{v}\in\{y|{\bm{A}}[:,y]={\mathbb{I}}\},$ (3)
@@
where ${\mathbb{L}}$ and ${\mathbb{I}}$ denote the language instruction and image tokens set, respectively.
Next, we average scores of all instruction tokens to obtain the estimate $\bar{{\bm{P}}}_{j}$ for $j$th vision token as
$\bar{{\bm{P}}}_{j}=\frac{1}{L_{t}}\sum\nolimits_{i=1}^{L_{t}}{\bm{P}}[i,j],\ \ j\in\{1,2,\ldots,L_{v}\},$ (4)
@@
where we use $\bar{{\bm{P}}}_{j}$ as the significance indicator for sparsification and a larger value in $\bar{{\bm{P}}}_{j}$ means higher significance for the corresponding token. Calculation of equation 4 costs $L_{t}\times L_{v}$ FLOPs, while the correlation matrix ${\bm{A}}$ and the indexing process is free, which benefits the inference efficiency.
@@
Relevant Text Token Selection. It is not appropriate to use all text tokens as a reference for visual sparsification. Figure 3 shows four representative cases where we compute the correlation between the prompt and the image. Case $3$ highlights Tylenol, Advil, ibuprofen, while top, sticker, fridge in case $4$ are significant, where a large proportion of question tokens in light red include little visual relevance. Therefore, it is unreasonable to make insignificant text tokens to rate vision tokens, and we need to select relevant text tokens (i.e., “raters”) for guidance.
@@
${\bm{S}}=\{\ i\ |\ {\bm{R}}[i]\geq m,\ \ i\in\{1,2,\ldots,L_{t}\}\}.$ (6)
${\bm{R}}=\frac{1}{L_{v}}\sum_{j=1}^{L_{v}}\text{Softmax}({\bm{H}}_{v}{{\bm{H}}_{q}}^{T}),$ (7)
where $m=\text{Mean}({\bm{R}})$ and only candidates who exceed the $m$ will become raters. The strategy ${\bm{S}}$ defines the index of selected raters from the candidate list containing $L_{t}$ tokens with priority. The equation 7 costs $L_{t}\times L_{v}\times 2D$ FLOPs, only computed once before entering the decoder.
Figure 3: Sample prompts from four representative multimodal benchmarks. The darker the word, the greater its relationship to the image and the more valuable it is for reference. We see that some words are irrelevant to the vision domain (e.g., prepositions and pronouns) and should not be considered for visual sparsification. Best viewed in color.
@@
$N=\lambda\times(L_{v}-\text{Rank}({\bm{P}})).$ (8)
We then remove $N$ visual tokens with the smallest values in ${\bm{P}}$. Notably, if the result of $N$ in a decoder layer is 00, we skip the layer and abandon sparsification. This stage requires $L_{t}\times L_{v}\times\text{min}(L_{t},L_{v})$ FLOPs for rank computation per layer.

### v3

@@ Abstract
In vision-language models (VLMs), visual tokens usually bear a significant amount of computational overhead despite sparsity of information in them when compared to text tokens. To address this, most existing methods learn a network to prune redundant visual tokens using certain training data. Differently, we propose a text-guided training-free token optimization mechanism dubbed SparseVLM that eliminates the need of extra parameters or fine-tuning costs. Given that visual tokens complement text tokens in VLM’s linguistic reasoning, we select relevant text tokens to rate the significance of visual tokens using self-attention matrices and, then, prune visual tokens using the proposed strategy to maximize sparsity while retaining information. In particular, we introduce a rank-based strategy to adaptively determine the sparsification ratio for each layer, alongside a token recycling method that compresses pruned tokens into more compact representations. Experimental results show that SparseVLM increases the efficiency of various VLMs in a number of image and video understanding tasks. For example, LLaVA when equipped with SparseVLM achieves 54% reduction in FLOPs, 37% decrease in CUDA latency while maintaining 97% of its original accuracy. Our code is available at https://github.com/Gumpest/SparseVLMs.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Comparison of visual token sparsification methods. Unlike previous methods with text-agnostic visual sparsification (c) e.g., VocoLLaMA (Ye et al., 2024), our SparseVLM (b) is guided by question prompts to select relevant visual patches from the image (a).
Benefiting from advancements in large language models (LLMs) (Radford et al., 2019; Brown et al., 2020; Achiam et al., 2023; Touvron et al., 2023; Peng et al., 2023; Bi et al., 2024), the realm of vision-language models (VLMs) has undergone significant progress. To combine visual signals with textual semantics, the mainstream practice in VLMs (Team et al., 2023; Bai et al., 2023; Chen et al., 2023; Li et al., 2024c, 2023a) employs sequential visual representation, where images are extracted into visual tokens and sent into an LLM decoder. With modal alignment and instruction fine-tuning (Du et al., 2021; Liu et al., 2023a; Zhu et al., 2023b), recent VLMs successfully adapt LLMs to the vision domain and inherit their perception and reasoning abilities.
Despite the promising performance, further incorporation of visual tokens inevitably introduces a huge memory and computational overhead when compared to LLMs, particularly for high-resolution images (Li et al., 2024c) and long videos (Lin et al., 2023). For instance, a $672\times 672$ image in LLaVA (Liu et al., 2024) yields $2304$ visual tokens that span over half of the context length. However, the information in images is typically more sparse than in natural languages (Marr, 2010), resulting in inefficiency when naïvely processing both modalities. To address this, existing methods extract more compact image representations by modifying the image encoder or projector (Alayrac et al., 2022; Li et al., 2024b; Dai et al., 2023; Cha et al., 2024). While some recent works further sparsify visual tokens during the decoding (Ye et al., 2024; Chen et al., 2024; Cai et al., 2024), they still ignore the guidance from the language tokens, which contradicts the multimodality paradigm. We argue that visual tokens should be sparsified adaptively based on the question prompt, as the model might focus on different parts (e.g., foreground or background) when dealing with various questions as shown in Figure 1. Furthermore, current approaches generally train a network to prune redundant visual tokens and require additional training data (Li et al., 2024b; Ye et al., 2024).
In this paper, we introduce a text-guided training-free framework dubbed SparseVLM for efficient vision language model inference. We reuse the self-attention matrix of visual-text tokens directly from the decoder layers without extra training parameters for sparsification. We ascertain that not all prompt tokens should be considered as some could be less relevant, which leads to inaccurate correlation results and downgrades the performance of sparse inference. Specifically, our SparseVLM first identifies text tokens strongly correlated with visual signals via cross-attention. Then, we measure the contribution of visual tokens to the selected visual-relevant text tokens (i.e., “raters”) and adaptively prune the insignificant visual tokens. Instead of directly discarding the pruned tokens, we further recycle and cluster them to reconstruct more compact tokens to minimize the loss of information. Due to the information density varying for different image inputs, we employ the rank of the attention matrix to indicate the redundancy level and set an adaptive sparsification ratio accordingly.
The proposed method is simple yet practical. It can act as a plug-and-play module to improve the efficiency of VLMs without additional fine-tuning. Extensive experiments demonstrate that our SparseVLM effectively reduces computational overhead of various VLMs without sacrificing their performance in a wide range of image and video understanding tasks. For instance, LLaVA (Liu et al., 2024) when armed with SparseVLM achieves a 4.5×4.5$\times$4.5 × compression rate while maintaining $97\%$ of its original performance. Alternatively, the CUDA latency can decrease by $37\%$ with only a $0.9\%$ drop in accuracy. To investigate the effectiveness of our method in video tasks, we further apply SparseVLM to VideoLLaVA (Lin et al., 2023) to compress frames with temporal dimension. Without complex design changes, SparseVLM can sparsify video frames into an adaptive number of visual tokens and outperform existing methods in video question-answering benchmarks. Our approach consistently outperforms prior state-of-the-art FastV method (Chen et al., 2024) by $11.2-17.3\%$ on LLaVA, $9.2-20.4\%$ on MiniGemini, and $34.4\%$ on VideoLLaVA when both have similar latencies.
Our main contributions are summarized as follows:
  • We introduce a novel sparsification framework dubbed SparseVLM. To the best of our knowledge, it is the first training-free approach that explores text-aware guidance for efficient VLM inference.
  • Particularly, we propose a strategy to select relevant text tokens as raters of visual tokens, a method to assess the significance of visual tokens followed by pruning of redundant visual tokens with a recycling mechanism to minimize the loss of information.
  • When applied to a number of VLMs, SparseVLM consistently outperforms prior state-of-the-art methods in various image and video understanding benchmarks.
@@ 2 Related Work
Vision-Language Models. Recent works on vision-language models (Liu et al., 2023a; Chen et al., 2023; Li et al., 2024c) improve multimodal comprehension and generation by processing longer visual token sequences. Moreover, the usage of higher-resolution images inevitably entails an exponential growth in the length of visual sequences. For example, LLaVA typically encodes $336\times 336$ images into $576$ tokens (Liu et al., 2024) with up to $672\times 672$ maximum resolution using $2880$ token sequences (Liu et al., 2023a). Similarly, mini-Gemini-HD (Li et al., 2024c) converts $1536\times 1536$ high resolution and $672\times 672$ low resolution images into $2880$ visual tokens. Moreover, comprehending videos or multiple images leads to increased token allocations for visual signals. For instance, the VideoLLaVA (Lin et al., 2023) and VideoPoet (Kondratyuk et al., 2023) use thousands of tokens to encode multiple image frames. However, large number of visual tokens results in a computational bottleneck. Further research on sparsification is urged to further unleash VLM capabilities.
@@
Visual Compression for VLMs. Compression of visual tokens is necessary because, on the one hand, their quantity is usually tens to hundreds of times that of language tokens. On the other hand, visual signals are inherently more sparse in information when compared to texts that have been produced by humans (Marr, 2010). Past efforts to address the above problem can be categorized into two directions. The first one centers on the compression of a vision tower or an efficient projection of vision modality. For instance, LLaMA-VID (Li et al., 2024b) exploits the Q-Former with the context token while DeCo (Yao et al., 2024) employs an adaptive pooling to downsample the visual tokens at the patch level. Methods that belong to the second direction (Ye et al., 2024; Chen et al., 2024; Cai et al., 2024) go deeper into the text modality and sparsify visual tokens during the LLM decoding stage, but they still lack the guidance from the text tokens. In this paper, SparseVLM takes note of this limitation and improves performance upon it.
3 Method
@@
In this section, we present our SparseVLM for efficient VLM inference. We first review the attention mechanism in VLMs and then introduce the detailed strategies for our visual sparsification including visual significance estimation, relevant text token selection, and sparsification level adaptation. We further propose token recycling to reduce information loss and provide a theoretical analysis of computation savings. The pipeline is shown in Figure 2.
@@
Figure 2: The architecture of SparseVLM. In stage (a), text raters are pre-selected before entering the sparsification LLM. In stage (b), adaptive sparsification is performed on LLM layers, involving computing redundancy and the recycling of reconstructed tokens.
@@ 3.1 Preliminary: Attention in VLM Decoders
VLM decoders typically rely on the causal self-attention from the original transformer architecture (Vaswani et al., 2017) for token interactions. Without loss of generality, we describe the single-head attention below. Formally, the self-attention matrix with logits ${\bm{A}}\in\mathbb{R}^{L\times L}$, where $L$ denotes the length of a sequence with all kinds of tokens e.g. text and visual, is computed by
@@
${\bm{A}}=\text{Attention}({\bm{Q}},{\bm{K}})=\textrm{Softmax}\left(\frac{{\bm{Q}}{\bm{K}}^{T}}{\sqrt{D}}\right),$ (1)
@@
where the scalar $D$ represents the matrix dimension, and the ${\bm{Q}}\in\mathbb{R}^{L\times D}$ and ${\bm{K}}\in\mathbb{R}^{L\times D}$ are the query and key matrices, respectively. The keys and queries in a self-attention layer are computed in parallel by using multi-layer perceptrons to transform the input hidden states ${\bm{H}}$ into a common space, where aligned interactions between modalities occur.
Often, the matrix ${\bm{A}}$ cannot be directly accessed due to FlashAttention-type (Dao et al., 2022) optimizations. Therefore, we develop an approach to extract ${\bm{A}}$ while maintaining compatibility with the FlashAttention when applying our sparsification. Please refer to the Appendix A.2.
3.2 Sparsification Guidance from Text to Vision
@@
Estimation of Visual Token Significance. For a multimodal model, we aim to estimate an impact of deleting a single token from one modality to other modalities. In the VLM case, we need to quantify how relevant a visual token is to text tokens in order to determine whether it can be pruned. Therefore, we naturally reuse the self-attention logits from VLM’s transformer layers as a reference since they already contain language-to-vision query results.
@@
In particular, we take the interaction between the query-dimensional part of the language modality and the key-dimensional part of the vision modality as the basis for sparsification priority matrix ${\bm{P}}\in\mathbb{R}^{L_{t}\times L_{v}}$, where $L_{t}$ and $L_{v}$ are the lengths of text and visual tokens, defined by
${\bm{P}}={\bm{A}}_{i,j},~{}\textrm{and}~{}(i,j)\in\{{\mathbb{L}},{\mathbb{I}}\},$ (2)
@@
where ${\mathbb{L}}$ and ${\mathbb{I}}$ denote the language instruction and image token sets, respectively.
Next, we obtain a vector $\tilde{{\bm{p}}}$ that estimates the significance of all visual tokens w.r.t. the text dimension as
$\tilde{{\bm{p}}}=\left[\tilde{p}_{1},\tilde{p}_{2},\ldots\tilde{p}_{L_{v}}\right]=\frac{1}{L_{t}}\sum\nolimits_{i=1}^{L_{t}}{\bm{P}}_{i},$ (3)
@@
where we use $\tilde{{\bm{p}}}$ as an indicator for sparsification and a larger value in $\tilde{{\bm{p}}}$ means higher significance of the corresponding visual token. Calculation of (3) costs $L_{t}\times L_{v}$ FLOPs only while the access to already computed ${\bm{A}}$ is considered as free, which is highlights low complexity of the SparseVLM.
@@
Relevant Text Token Selection. It is not appropriate to use all text tokens as a reference for visual sparsification. Figure 3 shows four representative cases where we compute the correlation between the prompt and the image. Case $3$ highlights Tylenol, Advil, ibuprofen, while sticker, fridge in case $4$ are significant, where a large proportion of question tokens in light red include little visual relevance. Therefore, it is unreasonable to make insignificant text tokens to rate visual tokens, and we need to select relevant text tokens (i.e., “raters”) for guidance.
@@
$\bm{s}=\{\ i\ |\ {\bm{r}}_{i}\geq m\},\ \ i\in\{1,2,\ldots,L_{t}\},$ (5)
${\bm{r}}=\frac{1}{L_{v}}\sum_{j=1}^{L_{v}}\left(\textrm{Softmax}\left({\bm{H}}_{v}{{\bm{H}}_{q}}^{T}\right)\right)_{j},$ (6)
where $m=\textrm{mean}({\bm{r}})$ and only candidates that exceed the $m$ threshold become raters. The strategy $\bm{s}$ contains the indices of selected raters from the candidate list of $L_{t}$ tokens. The (6) costs $L_{t}\times L_{v}\times 2D$ FLOPs that is only computed once before the decoder layer processing.
Figure 3: Sample prompts from four representative multimodal benchmarks. The darker the word, the greater its relationship to the image and the more valuable it is for reference. We see that some words are irrelevant to the vision domain (e.g., prepositions and pronouns) and should not be considered for visual sparsification. It is best viewed in color.
@@
$N=\lambda\times(L_{v}-\textrm{rank}({\bm{P}})).$ (7)
We then remove $N$ visual tokens with the smallest values in ${\bm{P}}$. Notably, if the result of $N$ in a decoder layer is 00, we skip the layer without sparsification. This stage requires $L_{t}\times L_{v}\times\text{min}(L_{t},L_{v})$ FLOPs for rank computation.